# Speculative Knowledge Distillation: Bridging the Teacher-Student Gap Through Interleaved Sampling

### v1

@@ Abstract
Recent advances in knowledge distillation (KD) have enabled smaller student models to approach the performance of larger teacher models. However, popular methods such as supervised KD and on-policy KD, are adversely impacted by the knowledge gaps between teacher-student in practical scenarios. Supervised KD suffers from a distribution mismatch between training with a static dataset and inference over final student-generated outputs. Conversely, on-policy KD, which uses student-generated samples for training, can suffer from low-quality training examples with which teacher models are not familiar, resulting in inaccurate teacher feedback. To address these limitations, we introduce Speculative Knowledge Distillation (SKD), a novel approach that leverages cooperation between student and teacher models to generate high-quality training data on-the-fly while aligning with the student’s inference-time distribution. In SKD, the student proposes tokens, and the teacher replaces poorly ranked ones based on its own distribution, transferring high-quality knowledge adaptively. We evaluate SKD on various text generation tasks, including translation, summarization, math, and instruction following, and show that SKD consistently outperforms existing KD methods across different domains, data sizes, and model initialization strategies.
@@ 4 Experimental Setup
Hyperparameters. For all fine-tuning process, we use learning rate $1$e-$5$, warmup ratio $0.1$ and $0.1$ drop out rate for all fine-tuning processes. All SFT checkpoint is trained for three epoches and we select the checkpoint with the lowest validation loss. Details about SFT training dataset can be found in Section 4.1. Without parameter search, we fixed acceptance criteria $K$ to be $25$ throughout main experiment and ablation ($K$=$25$ is a common choice for top-k sampling (Radford et al., 2019). See Appendix B for detailed study). We include training details and our hyper-parameter choice of baselines and SKD in the Appendix Section C.
@@
We evaluated SKD and baseline methods across three generation tasks: low-resource translation, dialogue summarization, and arithmetic reasoning. Task-specific KD often operates with training data around 1K samples (Quteineh et al., 2022). Thus, we randomly draw 1K samples from our task data, and further test low-data regime with 100 samples. Additionally, we performed experiments using 1K and 10K on a task-agnostic instruction-following task in the math domain. It’s important to note that different baseline KD methods have access to different training data. Supervised KD, SFT, and ImitKD have access to both $(x,y)$ pairs from the ground truth dataset, while on-policy and SKD only see the prompt $x$ and self-generate the target label during training. Since our assumption that the teacher model should be an expert in a specific task, we used full training data for each task to supervise the FT teacher model. For SFTed student initialization, we supervise the FT student model with our constructed training set $x,y$ (around 1000 instances). For reproducibility, all baselines and SKD use greedy decoding for evaluation. We include all results of supervised FT teacher model, supervised FT student model and instruction tuned student models in the Appendix D.
@@
We explored two types of knowledge distillation: task-specific (Quteineh et al., 2022) and task-agnostic (Agarwal et al., 2024). In task-specific distillation, training and inference data share the same format, and the task is known during training. In task-agnostic distillation, only the domain is known, while the specific task varies during deployment. We categorized distillation of translation, dialogue summarization, and arithmetic reasoning as task-specific and distillation of math instruction following as task-agnostic. For both math instruction following and arithmetic reasoning, we use chain-of-thought (Wei et al., 2023) prompts to generate answer. All prompt formats can be found in Appendix E. Details about parsing and evaluation for all tasks can be found in the Appendix F.
@@ 5 Results
We validate SKD’s effectiveness through comprehensive studies, including task-specific (Section 5.1) and task-agnostic distillation (Section 5.2), different model initializations (Section 4), low-data regimes (Section 5.4), and an ablation study on two-staged training (supervised KD followed by on-policy KD) (Section 5.5).
Table 1: SKD outperforms baseline KDs at in-domain tasks: low resource Assamese-to-English translation, dialogue summarization, arithmetic reasoning. We use Gemma-2b and Qwen-0.5B instruction tuned model (w/o SFT) as students and SFTed Gemma-7B-it and Qwen2-7B-it as teachers. $K$=$25$ for SKD.
@@ 5.1 Task-specific distillation
In this section, we investigated task-specific distillation, where the training and inference prompt-response sets share the same format, and the task is known during training. We study two model families, supervised FT Gemma-7B to instruction tuned Gemma-2B (Gemma-2B-it) and supervised FT Qwen2-7B to instruction tuned Qwen2-0.5B (Qwen2-0.5B-it). As shown in Table 1, the performance of supervised KD and on-policy KD is highly task-dependent, likely due to the model’s initial familiarity with each task (see Sec 4 for details). SKD, trained on Gemma-2B-it, consistently outperformed all baseline KD approaches across three in-domain text generation tasks.
@@ 5.2 Task-agnostic distillation
As illustrated in Figure 3, SKD initialized with Gemma-2B-it consistently outperformed supervised KD and on-policy KD across these four held-out sets when trained on 1k prompts. We also calculated the performance gain of SKD over SFT and reported baseline performance gains relative to SKD. Results showed that under 1k prompts, supervised KD and on-policy could approach close performance on simple tasks like SVAMP and ASDiv but still had a 20-40% performance gap compared to SKD (Second row on the left of Figure 3). The middle figure demonstrates SKD’s superiority over baselines on the Math task under 7 different concepts. Since the training set was task-agnostic instruction following data, it could be trained on a larger scale in practical scenarios. On the right side of Figure 3, we show that SKD can outperform or match all baseline methods with 10k prompts. Our results demonstrate SKD’s generalization capability under task-agnostic distillation. Detailed table result can also be found at Appendix D.
@@ 5.3 Different model initialization
In this section, we explore how different student model initialization impact the effectiveness of various KD approaches. As shown in Figure 4, on-policy KD initialized with an instruction-tuned student model can surprisingly degrade performance over time. This raises the question of whether a teacher’s token-wise probability distribution always provides beneficial guidance. Our findings suggest that if the initial student sample quality is insufficient, the student model may become trapped in a sub-optimal state, unable to effectively learn from the teacher’s supervision. On-policy KD’s low quality sample can be found in Sec 5.3 and Appendix G.
Table 2: SKD outperforms all baseline KDs with SFTed student initialization of instruction following model and domain-specific fine-tuned model across three text generation tasks. We use supervised SFTed Gemma-7B-it and Qwen2-7B-it model as teachers and supervised SFT Gemma-2B-it and Qwen2-0.5B-it as students. SKD is performed under $K$=$25$.
@@
Figure 4: Comparison between SKD and baseline KD methods with different model initializations. We find that on-policy methods struggles when the student model starts with a poor initialization, leading to performance degradation and becoming stuck at a low level throughout training. On-policy KD requires student model to have a good initialization. In contrast, SKD outperforms supervised and on-policy KD under both IT and supervised FT Gemma-2B initialization. On-policy KD’s low quality sample can be found in Appendix G. SKD is performed under $K$=$25$.
As shown in Table 2, SKD with SFT-initialized students consistently outperforms all baselines across both model families. While on-policy KD can narrow the gap with SKD through a two-stage training process (supervised fine-tuning + KD) for Gemma-2B students, it lags behind SKD for smaller models like Qwen-0.5B. Additionally, in Section 5.4 and Appendix Section H, we demonstrate that SFT can lead to over-fitting in student models under low-data regimes, resulting in sub-optimal performance in post-training KDs. Furthermore, despite some advantages over supervised and on-policy KD with IT-initialized student models, ImitKD does not exhibit clear benefits with higher quality initialized student models. This indicates that SKD’s token-level mixing of teacher and student model tokens can achieve more optimal performance than sequence-level mixing.
@@ 5.4 Low data regime
We further investigated SKD under an extremely low data setting with 100 data points per task (10% of the training data). As shown in Table 3, SKD outperformed supervised KD and on-policy KD for all three tasks when initialized with IT and SFTed Gemma-2B, demonstrating SKD’s superior performance in this extreme low-data setting. Notably, both supervised KD and SKD exhibited lower or comparable performance when switching from instruction-tuned to SFTed initialization. This is attributed to unavoidable overfitting during SFT with only 100 data points. In Appendix Section H, we illustrate how overfitting at the SFT stage can lead to suboptimal performance in post-training KDs. This finding suggests that end-to-end SKD, capable of training from any base model and bypassing the SFT stage, offers a more optimal solution than two-stage on-policy KD (SFT+KD) in extreme low-data scenarios.
@@ 5.5 Two-staged training
Given that SKD transitions from supervised KD-like behavior at the beginning of training to on-policy KD towards the end, we investigated whether a simple approach of training with supervised KD for the first half of iterations and self-generated samples for the second half could achieve SKD-like performance. We conducted an ablation study by training the student model with supervised KD first, followed by on-policy KD (see Appendix Section I for details). As shown in Figure 5, SKD significantly outperforms this mixed training strategy. Surprisingly, this heuristic baseline was often outperformed by either supervised KD or on-policy KD across three tasks and two model initializations. This is likely due to the data discrepancy encountered during the training process – naively mixing samples from the teacher and student in two different stages lacks natural confidence measurement of the tokens; therefore, simply combining supervised and on-policy KD can even degrade the performance of either approach. This underscores the importance of our adaptive SKD, which implicitly covers intermediate transitions between two model behaviors.
@@ 5.6 Summary and Additional results
SKD consistently outperforms baselines in both task-specific and task-agnostic distillation. Unlike on-policy methods, which are sensitive to student-generated sample quality, SKD maintains its advantage even with low-quality student model initialization. As shown in Appendix H, overfitting in SFT can hinder post-training KD’s performance. SKD’s end-to-end approach, bypassing SFT, provides a significant benefit, especially in low-data scenarios. Merging supervised and on-policy KD naively can lead to performance degradation due to data inconsistencies.
Speculative decoding, as described in (Leviathan et al., 2023), is a crucial technique for accelerating LLM inference. The quality of the draft model used in this process directly influences the speed-up achieved. As demonstrated in Appendix J, our SKD-trained draft model significantly outperforms the instruction-tuned base model in terms of token acceptance rate ($71$% to $85$% higher), leading to a $1.2$X speed-up in speculative decoding.
@@ 8 REPRODUCIBILITY STATEMENT
We discussed our model implementation, hyper-parameters, baseline details in Section 4. We include information about our dataset and evaluation metrics in Section 4.1. We discuss training and hyper-parameter details in Appendix C. We include examples of our prompt and response for all tasks in Appendix E. We include additional information about our evaluation metrics in Appendix F. All data, model and evaluation metrics are open sourced and available for research community. Upon the camera ready of the paper, we will include a link to a publicly accessible code repository, including all implementations about baseline KDs and SKD, running instructions and hyper-parameter settings. All results reported in this paper can be fully reproduced by the community.
@@
Appendix B Can SKD outperform baselines under different acceptance criteria?
@@
Appendix C Training and Hyperparameter Details for Baselines and SKD
@@
Appendix D Additional results
Tables 4 and 5 show the performance of Gemma-2B-it (student), Qwen-0.5B-it (students) and supervised FT teacher models (Gemma-7B-it and Qwen-7B-it) under task-specific and task-agnostic distillation. Table 6 presents task-agnostic instruction tuning results for baseline and SKD models across four held-out test sets, using both low-data (1k prompts) and full-data (10k prompts) training regimes. These results align with those shown in Figure 3.
Appendix E Task input and output Format

### v2

@@ Abstract
Recent advances in knowledge distillation (KD) have enabled smaller student models to approach the performance of larger teacher models. However, popular methods such as supervised KD and on-policy KD, are adversely impacted by the knowledge gaps between teacher-student in practical scenarios. Supervised KD suffers from a distribution mismatch between training with a static dataset and inference over final student-generated outputs. Conversely, on-policy KD, which uses student-generated samples for training, can suffer from low-quality training examples with which teacher models are not familiar, resulting in inaccurate teacher feedback. To address these limitations, we introduce Speculative Knowledge Distillation (SKD), a novel approach that leverages cooperation between student and teacher models to generate high-quality training data on-the-fly while aligning with the student’s inference-time distribution. In SKD, the student proposes tokens, and the teacher replaces poorly ranked ones based on its own distribution, transferring high-quality knowledge adaptively. We evaluate SKD on various text generation tasks, including translation, summarization, math, and instruction following, and show that SKD consistently outperforms existing KD methods across different domains, data sizes, and model initialization strategies. The code and data are released at https://github.com/xu1998hz/skd.
@@ 4 Experimental Setup
Hyperparameters. For all fine-tuning process, we use learning rate $1$e-$5$, warmup ratio $0.1$ and $0.1$ drop out rate for all fine-tuning processes. All SFT checkpoint is trained for three epoches and we select the checkpoint with the lowest validation loss. Details about SFT training dataset can be found in Section 4.1. Without parameter search, we fixed acceptance criteria $K$ to be $25$ throughout main experiment and ablation ($K$=$25$ is a common choice for top-k sampling (Radford et al., 2019). See Appendix B for detailed study). We include training details and our hyper-parameter choice of baselines and SKD in the Appendix Section D.
@@
We evaluated SKD and baseline methods across three generation tasks: low-resource translation, dialogue summarization, and arithmetic reasoning. Task-specific KD often operates with training data around 1K samples (Quteineh et al., 2022). Thus, we randomly draw 1K samples from our task data, and further test low-data regime with 100 samples. Additionally, we performed experiments using 1K and 10K on a task-agnostic instruction-following task in the math domain. It’s important to note that different baseline KD methods have access to different training data. Supervised KD, SFT, and ImitKD have access to both $(x,y)$ pairs from the ground truth dataset, while on-policy and SKD only see the prompt $x$ and self-generate the target label during training. Since our assumption that the teacher model should be an expert in a specific task, we used full training data for each task to supervise the FT teacher model. For SFTed student initialization, we supervise the FT student model with our constructed training set $x,y$ (around 1000 instances). For reproducibility, all baselines and SKD use greedy decoding for evaluation. We include all results of supervised FT teacher model, supervised FT student model and instruction tuned student models in the Appendix E.
@@
We explored two types of knowledge distillation: task-specific (Quteineh et al., 2022) and task-agnostic (Agarwal et al., 2024). In task-specific distillation, training and inference data share the same format, and the task is known during training. In task-agnostic distillation, only the domain is known, while the specific task varies during deployment. We categorized distillation of translation, dialogue summarization, and arithmetic reasoning as task-specific and distillation of math instruction following as task-agnostic. For both math instruction following and arithmetic reasoning, we use chain-of-thought (Wei et al., 2023) prompts to generate answer. All prompt formats can be found in Appendix F. Details about parsing and evaluation for all tasks can be found in the Appendix G.
@@ 5 Results
We validate SKD’s effectiveness through comprehensive studies, including task-specific (Section 5.1) and task-agnostic distillation (Section 5.2), different model initializations (Section 5.3), low-data regimes (Section 5.4), and an ablation study on two-staged training (supervised KD followed by on-policy KD) (Section 5.5).
Table 1: SKD outperforms baseline KDs at in-domain tasks: low resource Assamese-to-English translation, dialogue summarization, arithmetic reasoning. We use Gemma-2b and Qwen-0.5B instruction tuned model (w/o SFT) as students and SFTed Gemma-7B-it and Qwen2-7B-it as teachers. $K$=$25$ for SKD. We conducted a permutation significance test for SKD against all baselines for translation and summarization. SKD outperforms all baselines (p$<$0.05) significantly except for SKD vs ImitKD at Gemma-7B to Gemma-2B-IT on summarization and SKD vs supervised KD/SFT at Qwen-7B to Qwen-0.5B-it on translation.
@@ 5.1 Task-specific distillation
In this section, we investigated task-specific distillation, where the training and inference prompt-response sets share the same format, and the task is known during training. We study two model families, supervised FT Gemma-7B to instruction tuned Gemma-2B (Gemma-2B-it) and supervised FT Qwen2-7B to instruction tuned Qwen2-0.5B (Qwen2-0.5B-it). As shown in Table 1, the performance of supervised KD and on-policy KD is highly task-dependent, likely due to the model’s initial familiarity with each task (see Sec 5.3 for details). SKD, trained on Gemma-2B-it, consistently outperformed all baseline KD approaches across three in-domain text generation tasks.
@@ 5.2 Task-agnostic distillation
As illustrated in Figure 3, SKD initialized with Gemma-2B-it consistently outperformed supervised KD and on-policy KD across these four held-out sets when trained on 1k prompts. We also calculated the performance gain of SKD over SFT and reported baseline performance gains relative to SKD. Results showed that under 1k prompts, supervised KD and on-policy could approach close performance on simple tasks like SVAMP and ASDiv but still had a 20-40% performance gap compared to SKD (Second row on the left of Figure 3). The middle figure demonstrates SKD’s superiority over baselines on the Math task under 7 different concepts. Since the training set was task-agnostic instruction following data, it could be trained on a larger scale in practical scenarios. On the right side of Figure 3, we show that SKD can outperform or match all baseline methods with 10k prompts. Our results demonstrate SKD’s generalization capability under task-agnostic distillation. Detailed table result can also be found at Appendix E.
@@ 5.3 Different model initialization
In this section, we explore how different student model initialization impact the effectiveness of various KD approaches. As shown in Figure 4, on-policy KD initialized with an instruction-tuned student model can surprisingly degrade performance over time. This raises the question of whether a teacher’s token-wise probability distribution always provides beneficial guidance. Our findings suggest that if the initial student sample quality is insufficient, the student model may become trapped in a sub-optimal state, unable to effectively learn from the teacher’s supervision. On-policy KD’s low quality sample can be found in Sec 5.3 and Appendix H.
Table 2: SKD outperforms all baseline KDs with SFTed student initialization of instruction following model and domain-specific fine-tuned model across three text generation tasks. We use supervised SFTed Gemma-7B-it and Qwen2-7B-it model as teachers and supervised SFT Gemma-2B-it and Qwen2-0.5B-it as students. SKD is performed under $K$=$25$. We conducted a permutation significance test for SKD against all baselines for translation and summarization. SKD outperforms all baselines (p$<$0.05) significantly.
@@
Figure 4: Comparison between SKD and baseline KD methods with different model initializations. We find that on-policy methods struggles when the student model starts with a poor initialization, leading to performance degradation and becoming stuck at a low level throughout training. On-policy KD requires student model to have a good initialization. In contrast, SKD outperforms supervised and on-policy KD under both IT and supervised FT Gemma-2B initialization. On-policy KD’s low quality sample can be found in Appendix H. SKD is performed under $K$=$25$.
As shown in Table 2, SKD with SFT-initialized students consistently outperforms all baselines across both model families. While on-policy KD can narrow the gap with SKD through a two-stage training process (supervised fine-tuning + KD) for Gemma-2B students, it lags behind SKD for smaller models like Qwen-0.5B. Additionally, in Section 5.4 and Appendix Section I, we demonstrate that SFT can lead to over-fitting in student models under low-data regimes, resulting in sub-optimal performance in post-training KDs. Furthermore, despite some advantages over supervised and on-policy KD with IT-initialized student models, ImitKD does not exhibit clear benefits with higher quality initialized student models. This indicates that SKD’s token-level mixing of teacher and student model tokens can achieve more optimal performance than sequence-level mixing.
@@ 5.4 Low data regime
We further investigated SKD under an extremely low data setting with 100 data points per task (10% of the training data). As shown in Table 3, SKD outperformed supervised KD and on-policy KD for all three tasks when initialized with IT and SFTed Gemma-2B, demonstrating SKD’s superior performance in this extreme low-data setting. Notably, both supervised KD and SKD exhibited lower or comparable performance when switching from instruction-tuned to SFTed initialization. This is attributed to unavoidable overfitting during SFT with only 100 data points. In Appendix Section I, we illustrate how overfitting at the SFT stage can lead to suboptimal performance in post-training KDs. This finding suggests that end-to-end SKD, capable of training from any base model and bypassing the SFT stage, offers a more optimal solution than two-stage on-policy KD (SFT+KD) in extreme low-data scenarios.
@@ 5.5 Two-staged training
Given that SKD transitions from supervised KD-like behavior at the beginning of training to on-policy KD towards the end, we investigated whether a simple approach of training with supervised KD for the first half of iterations and self-generated samples for the second half could achieve SKD-like performance. We conducted an ablation study by training the student model with supervised KD first, followed by on-policy KD (see Appendix Section J for details). As shown in Figure 5, SKD significantly outperforms this mixed training strategy. Surprisingly, this heuristic baseline was often outperformed by either supervised KD or on-policy KD across three tasks and two model initializations. This is likely due to the data discrepancy encountered during the training process – naively mixing samples from the teacher and student in two different stages lacks natural confidence measurement of the tokens; therefore, simply combining supervised and on-policy KD can even degrade the performance of either approach. This underscores the importance of our adaptive SKD, which implicitly covers intermediate transitions between two model behaviors.
@@ 5.6 Summary and Additional results
SKD consistently outperforms baselines in both task-specific and task-agnostic distillation. Unlike on-policy methods, which are sensitive to student-generated sample quality, SKD maintains its advantage even with low-quality student model initialization.
In Appendix B, we demonstrated that a wide range of KK values can outperform both supervised and on-policy KD, and that optimal KK values can be further explored across tasks. Additionally, Appendix C presents a preliminary experiment on an adaptive KK criterion, highlighting a promising direction for future research. Furthermore, as discussed in Appendix I, overfitting in SFT can impede the performance of post-training KD. By adopting an end-to-end approach that bypasses SFT, SKD offers a significant advantage, particularly in low-data scenarios.
Speculative decoding, as described in (Leviathan et al., 2023), is a crucial technique for accelerating LLM inference. The quality of the draft model used in this process directly influences the speed-up achieved. As demonstrated in Appendix K, our SKD-trained draft model significantly outperforms the instruction-tuned base model in terms of token acceptance rate ($71$% to $85$% higher), leading to a $1.2$X speed-up in speculative decoding.
In Appendices L and M, we analyze the rejection rate of student tokens and estimate the theoretical computational cost of SKD, finding that SKD reduces computation by 50% compared to directly sampling from the teacher model. Additionally, in Appendix N, SKD demonstrates superior validation convergence over baseline methods, achieving the lowest validation loss across tasks and initializations. Finally, in Appendix O, we quantify sample OOD issues across different KD approaches.
@@ 8 REPRODUCIBILITY STATEMENT
We discussed our model implementation, hyper-parameters, baseline details in Section 4. We include information about our dataset and evaluation metrics in Section 4.1. We discuss training and hyper-parameter details in Appendix D. We include examples of our prompt and response for all tasks in Appendix F. We include additional information about our evaluation metrics in Appendix G. All data, model and evaluation metrics are open sourced and available for research community. Upon the camera ready of the paper, we will include a link to a publicly accessible code repository, including all implementations about baseline KDs and SKD, running instructions and hyper-parameter settings. All results reported in this paper can be fully reproduced by the community.
@@
Appendix B Can SKD outperform baselines under different $K$ choices?
@@
Appendix C Adaptive K study
We explored experiments by adjusting $K$ during training steps (decreasing $K$). However, we didn’t obtain positive signals (it didn’t beat constant $K$ approach). Our hypothesis is that SKD assumes that as the model improves, it naturally acts like on-policy KD. However, heuristically decreasing $K$ can force it to replace the token and behave like supervised KD, which introduces off-policy tokens, hurting the performance of the model. However, we encourage future works to explore more into this adaptive $K$ direction.
We also conducted experiments using adaptive $K$ strategy at decoding steps. Due to the autoregressive nature of our base models, token generation becomes more deterministic with longer sequences. This raises the question of whether reducing $K$ as the sequence grows could further improve performance. We experimented with a linear decay of $K$, decreasing it by for each additional generated token. We set a minimum $K$ to ensure our algorithm does not degenerate to supervised KD towards the end of decoding. In Table 4, we showed that adaptive SKD outperforms most of KD baselines across translation and summarization, it does not outperform constant SKD with K=25.
Table 4: In this Table, we showed that performance under adaptive $K$. We initialized $K$ with $50$ and $25$ respectively and $K$ is reduced linearly with increasing decoding step. We also setup clipping $K$ value so that adaptive SKD pipeline will not degenerate to supervised KD at later decoding stage. Cliiping value is set to be $25$ for $k$ is initialized at $50$ and $15$ for $k$ is initialized at $25$. Although adaptive SKD outperforms most of KD baselines across translation and summarization, it does not outperform constant SKD with $K$=$25$.
Translation (Measured by COMET) Summarization (Measured by ROUGE-L)
$K$ from $50$ to $25$ $K$ from $25$ to $15$ $K$=$25$ $K$ from $50$ to $25$ $K$ from $25$ to $15$ $K$=$25$
Performance 75.1 75.0 75.3 25.2 28.7 29.1
We also encourage future work to study Top-$P$ criteria (Holtzman et al., 2020) for SKD pipeline. We believe that combing Top-$P$ and Top-$K$ criteria can lead to better performance of SKD pipeline.
Appendix D Training and Hyperparameter Details for Baselines and SKD
@@
Appendix E Additional results
Tables 5 and 6 show the performance of Gemma-2B-it (student), Qwen-0.5B-it (students) and supervised FT teacher models (Gemma-7B-it and Qwen-7B-it) under task-specific and task-agnostic distillation. Table 7 presents task-agnostic instruction tuning results for baseline and SKD models across four held-out test sets, using both low-data (1k prompts) and full-data (10k prompts) training regimes. These results align with those shown in Figure 3.
Appendix F Task input and output Format